Circular No. 6195

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995T IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of V ∼ 18.0 at α = 22^{h}27^{m}12^{s}.41, δ = −9^{o}29′44′′.4 (equinox 2000.0), 6′′.1 west and 2′′.3 south of the center of an anonymous galaxy of m_{pg} ∼ 17.4. The object was found on tech pan films obtained by D. Albanese and himself on July 25.09 and 27.05 UT with the 0.9-m Schmidt. Nothing is visible at the location on the Palomar Sky Survey prints down to B = 21 and R = 20. A. V. Filippenko, D. C. Leonard and A. J. Barth, University of California at Berkeley, report that preliminary inspection of uncalibrated CCD spectra (range 310–800 nm, resolution 0.7 nm) obtained on Aug. 1 with the 3-m Shane reflector at the Lick Observatory suggests that the object is a type Ia supernova, probably ∼ 1–2 weeks past maximum brightness. The redshift of the parent galaxy, determined from strong emission lines, is 0.056.

WGA J1047.1+6335 AND J1802.1+1804 K. P. Singh, P. Barrett, E. Schlegel and N. E. White, Goddard Space

Flight Center; P. Szkody, A. Silber, D. W. Hoard and E. Fierce, University of Washington, report: “We have found two new cataclysmic variables based on their x-ray and optical characteristics. The sources were discovered during a search for the ultrasoft sources in the ROSAT PSPC pointed observations, data from which have been archived at HEASARC and are publicly available as WGACAT (cf. IAUC 6100). The x-ray spectrum of WGA J1047.1+6335 shows a dominant blackbody with kT = 40–50 eV and unabsorbed flux of 3.5 × 10^{−13} erg cm^{−2} s^{−1} in the 0.1–0.5-keV energy range; no x-rays > 0.5 keV are detected. The optical counterpart at α = 10^{h}47^{m}09^{s}.9, δ = +63^{o}35′13′′ (equinox 2000.0) has mag ∼ 19.4. Strong emission lines of H and He II are present, and a period of ∼ 40 min is seen in photometry. This object probably belongs to a new class of soft intermediate polars (Haberl and Motch 1995, A.Ap. 297, L37). An x-ray spectrum of WGA J1802.1+1804 also shows a dominant soft blackbody with kT ∼ 30 eV and a bremsstrahlung component with kT > 1 keV, the blackbody flux being about 10^{3} times the flux in the bremsstrahlung component. The total unabsorbed x-ray flux at 0.1–2.0 keV is 4.1 × 10^{−9} erg cm^{−2} s^{−1}. The optical counterpart is located at α = 18^{h}02^{m}06^{s}.5, δ = +18^{o}04′43′′ (equinox 2000.0). Of mag 15, it has strong emission lines in its optical spectrum, with He II 469 nm stronger than Hβ. It is likely to be an AM Her-type object. An optical and x-ray period of ∼ 113 min has also been detected. Follow-up polarimetric observations of both objects are encouraged.”

1995 August 1 Brian G. Marsden